OpenStudy (anonymous):

The base of a rectangle is 22 centimeters and the height is 4 centimeters a similar rectangle has a height of 6 centimeters what is the base of the second rectangle

OpenStudy (zpupster):

ratio 4/22=6/x or 4x=6*22 sole for x